OpenStudy (anonymous):

A cold beer initially at 33°F warms up to 38°F in 3 min while sitting in a room of temperature 70°F. How warm will the beer be if left out for 18 min?

OpenStudy (anonymous):

using Malthusian law, i keep getting 76.9 but its wrong -.-

OpenStudy (anonymous):

ok the initial difference in temperature is \[70-33=37\] after 3 minutes it is \[70-38=32\] degrees so you can use \[37\times (\frac{32}{37})^{\frac{t}{3}}\] for the temperature difference after t minutes

OpenStudy (anonymous):

try \[37(\frac{32}{37})^{6}\] and see what you get.

OpenStudy (anonymous):

ok ill check it out right now, thanks alot

OpenStudy (anonymous):

i get about 15.54 as the differnce between the room and the beer, making the beer 70 - 15.54 = 54.46 rounded

OpenStudy (anonymous):

yea thats the answer, im going to walk my self through the logic. Thanks for the info